OpenStudy (anonymous):

evaluate the expression: 2x(6+3)x9=?

OpenStudy (anonymous):

2

OpenStudy (idealist10):

2*(6+3)*9=2*9*9=18*9=162 Answer: 162

OpenStudy (anonymous):

no its 2

OpenStudy (idealist10):

Can you show your work?

OpenStudy (anonymous):

me?

OpenStudy (anonymous):

2 x 6 =12 +3=15 x 9=2

OpenStudy (anonymous):

Thanks! @Idealist10

OpenStudy (idealist10):

You're welcome! :)

OpenStudy (anonymous):

its 2!!!